OpenStudy (lena772):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). y = 2x + 6 y = -one halfx + 3 y = -one halfx + 6 y = 2x + 3

ganeshie8 (ganeshie8):

wats the slope of given line y = 2x+2 ?

ganeshie8 (ganeshie8):

from that, u can find the slope of perpendicular line

OpenStudy (lena772):

2

ganeshie8 (ganeshie8):

yes, wats the slope of perpendicular lien ?

OpenStudy (lena772):

-1/2

OpenStudy (lena772):

because the slope of a perpendicular line is the negative recipricol of the slope of the parallel line

OpenStudy (lena772):

i think

OpenStudy (lena772):

3=-1/2*6+b 6=b

OpenStudy (lena772):

y=-1/2x+6 right?

ganeshie8 (ganeshie8):

Correct ! good job !!